# What Matters in Data for DPO?

### v1

@@ B.1 Dataset Details
Some may doubt the reliance on reward model (RM) scores as quality criteria, given known calibration limitations that prevent these scores from being perfect human preference estimators. We justify this design through two principled arguments. First, our experiments require aggregating responses from heterogeneous sources, where a unified quantitative metric becomes indispensable for quality-aware sample reorganization. Second, empirical evidence from our LLM-as-a-judge comparison validates RM’s practical superiority: on the Tulu3 UltraFeedback dataset, RM and LLM evaluators disagreed on response quality rankings for 10,000 prompts. Crucially, when training DPO models on datasets filtered by each method, RM-based selection achieved better performance, demonstrating its operational robustness. Recent work has also revealed the possible limitation of LLM-as-a-judge methods (Li et al., 2025). Although we do not claim RM scores are intrinsically perfect, their empirical stability and cross-domain applicability make them a functionally optimal choice for our multi-source quality stratification objectives.
@@ B.2 On-policy Data Generation
In the context of Direct Preference Optimization (DPO), on-policy data refers to preference pairs generated using the policy model that is currently being trained or a recent checkpoint thereof. Training with such data allows the model to learn from its own evolving capabilities. Several methods leverage on-policy data, including fully online approaches like online DPO (Guo et al., 2024) and online iterative DPO (Dong et al., 2024a). While potentially effective, these online methods often require frequent interaction with a preference judge (e.g., a human annotator or a reward model) during the training loop, which can significantly increase computational and annotation costs.
An alternative strategy, which balances the benefits of on-policy data with practical constraints, involves generating a batch of on-policy data offline before commencing or resuming DPO training. This generated data can then be mixed with existing off-policy datasets (Lambert et al., 2024; Deng et al., 2025). In this paper, we adopt a similar offline generation approach, closely following the methodology described by Lambert et al. (2024). The specific process for generating our on-policy preference data is detailed in Algorithm 1.
@@ B.4 Evaluation
We leverage the Tulu3 evaluation pipeline except for AlpacaEval-2 to exclude the possible evaluation data leakage. For the AlpacaEval-2 assessment, we adopt a generation config of beam-search multinomial sampling with num_beams=3 and temperature=1.0. For the other benchmarks, we use the default configuration as described in Lambert et al. (2024). The score metrics used in our experiments are presented in Table 7.

### v3

@@ B.1 Dataset Details
Some may doubt the reliance on reward model (RM) scores as quality criteria, given known calibration limitations that prevent these scores from being perfect human preference estimators. We justify this design through two principled arguments. First, our experiments require aggregating responses from heterogeneous sources, where a unified quantitative metric becomes indispensable for quality-aware sample reorganization. Second, empirical evidence from our LLM-as-a-judge comparison validates RM’s practical superiority: on the Tulu3 UltraFeedback dataset, RM and LLM evaluators disagreed on response quality rankings for 10,000 prompts. Crucially, when training DPO models on datasets filtered by each method, RM-based selection achieved better performance, demonstrating its operational robustness. Recent work has also revealed the possible limitation of LLM-as-a-judge methods (li2025preference). Although we do not claim RM scores are intrinsically perfect, their empirical stability and cross-domain applicability make them a functionally optimal choice for our multi-source quality stratification objectives.
@@ B.2 On-policy Data Generation
In the context of Direct Preference Optimization (DPO), on-policy data refers to preference pairs generated using the policy model that is currently being trained or a recent checkpoint thereof. Training with such data allows the model to learn from its own evolving capabilities. Several methods leverage on-policy data, including fully online approaches like online DPO (guo2024direct) and online iterative DPO (dong2024rlhf). While potentially effective, these online methods often require frequent interaction with a preference judge (e.g., a human annotator or a reward model) during the training loop, which can significantly increase computational and annotation costs.
An alternative strategy, which balances the benefits of on-policy data with practical constraints, involves generating a batch of on-policy data offline before commencing or resuming DPO training. This generated data can then be mixed with existing off-policy datasets (lambert2024t; deng2025less). In this paper, we adopt a similar offline generation approach, closely following the methodology described by lambert2024t. The specific process for generating our on-policy preference data is detailed in Algorithm 1.
@@ B.4 Evaluation
We leverage the Tulu3 evaluation pipeline except for AlpacaEval-2 to exclude the possible evaluation data leakage. For the AlpacaEval-2 assessment, we adopt a generation config of beam-search multinomial sampling with num_beams=3 and temperature=1.0. For the other benchmarks, we use the default configuration as described in lambert2024t. The score metrics used in our experiments are presented in Table 7.